Circular No. 8087

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF JUPITER S. S. Sheppard, University of Hawaii (UH), reports the discovery of

seven new satellites of Jupiter, with the astrometry, orbital elements, and ephemerides given on MPEC 2003-E11. CCD observations were obtained during Feb. 5–Mar. 4 by Sheppard, D. C. Jewitt, J. Kleyna, Y. R. Fernandez, and H. H. Hsieh at Mauna Kea with the 8.3-m Subaru telescope, the 3.6-m Canada-France-Hawaii Telescope, and the 2.2-m UH telescope. The very preliminary orbital elements by B. G. Marsden indicate that two objects have direct orbits and five have retrograde orbits: S/2003 J 1, e = 0.79, i = 35^{o}, P = 236 days; S/2003 J 2, e = 0.38, i = 152^{o}, P = 983 days; S/2003 J 3, e = 0.24, i = 144^{o}, P = 505 days; S/2003 J 4, e = 0.20, i = 145^{o}, P = 722 days; S/2003 J 5, e = 0.21, i = 165^{o}, P = 761 days; S/2003 J 6, e = 0.76, i = 22^{o}, P = 234 days; S/2003 J 7, e = 0.41, i = 159^{o}, P = 747 days. Absolute magnitudes range from 15.6 to 16.9.

SUPERNOVA 2003bg IN MCG –05-10-15 A. M. Soderberg, S. R. Kulkarni, and E. Berger, California Institute

of Technology; and D. A. Frail, National Radio Astronomy Observatory, report that an observation of SN 2003bg with the Very Large Array at 8.46 GHz on Mar. 4.07 UT reveals a radio source with a flux density of 2.50 ± 0.06 mJy that is coincident with the supernova (IAUC 8082). Estimated as being ∼ 2 weeks past the explosion time (IAUC 8084), SN 2003bg has a radio luminosity that is 2 percent of that observed for SN 1998bw at 14 days after the supernova explosion (Kulkarni et al. 1998, Nature 395, 663).

GRB 030227 R. González-Riestra, M. Guainazzi, N. Loiseau, P. M. Rodríguez-

Pascual, M. Santos-Lleó, N. Schartel, B. Juárez, R. Pérez-Martínez, and M. Gilomo, XMM-Newton Science Operations Centre, report that an XMM-Newton target-of-opportunity observation of GRB 030227 on Feb. 27.7 UT detected a bright source at α = 4^{h}57^{m}33^{s}.3, δ = +20^{o}29′04′′ (equinox 2000.0; uncertainty ∼ 4′′) with the EPIC-pn and MOS cameras during a 49000-s exposure. The source faded (in the band 0.2–5 keV) from 0.15 to 0.07 EPIC-pn counts/s during the observation. The spectrum is consistent with a hydrogen column density of 3 × 10^{21} cm^{−2}. The observed flux in the range 2–10 keV is 5.5 × 10^{−13} erg cm^{−2} s^{−1}.

2003 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green